# Design2GarmentCode: Turning Design Concepts to Tangible Garments Through Program Synthesis

### v2

@@ Abstract
Sewing patterns, the essential blueprints for fabric cutting and tailoring, act as a crucial bridge between design concepts and producible garments. However, existing uni-modal sewing pattern generation models struggle to effectively encode complex design concepts with a multi-modal nature and correlate them with vectorized sewing patterns that possess precise geometric structures and intricate sewing relations. In this work, we propose a novel sewing pattern generation approach Design2GarmentCode based on Large Multimodal Models (LMMs), to generate parametric pattern-making programs from multi-modal design concepts. LMM offers an intuitive interface for interpreting diverse design inputs, while pattern-making programs could serve as well-structured and semantically meaningful representations of sewing patterns, and act as a robust bridge connecting the cross-domain pattern-making knowledge embedded in LMMs with vectorized sewing patterns. Experimental results demonstrate that our method can flexibly handle various complex design expressions such as images, textual descriptions, designer sketches, or their combinations, and convert them into size-precise sewing patterns with correct stitches. Compared to previous methods, our approach significantly enhances training efficiency, generation quality, and authoring flexibility. Our code and data will be publicly available.
@@ 2.2 LLMs for Program Synthesis
Recent advancements in program synthesis and code generation using large language models (LLMs) have laid essential groundwork for systems like Design2GarmentCode, which generate structured garment code from multi-modal design inputs. Earlier researches like Codex [11] and AlphaCode [35] demonstrated the effectiveness of LLMs in generating complex, task-specific code with high syntax accuracy, showcasing potential in scenarios requiring precise parametric coding. These models [42, 58] highlight how LLMs, when sufficiently trained, can transform natural language inputs into executable code, a capability directly relevant to generating garment codes that follow complex pattern-making syntax.
Additionally, researchers [54, 18, 9, 19, 62] have been exploring explore multi-modal models that integrate visual aids, such as flowcharts and UML diagrams, into LLM training to enhance models’ comprehension of complex structures and flow. These models particularly emphasize the need for semantic understanding and adaptability, which are critical in working with domain-specific languages (DSLs) like GarmentCode.
@@ 2.3 Neurosymbolic Models
Procedural/symbolic models and learned/neural models have complementary strengths and weaknesses. Neurosymbolic models [46] tend to combine the strengths of both paradigms and propose to generate visual data using symbolic programs augmented with AI/ML techniques. The neurosymbolic pipeline typically includes task specification, program synthesis using a DSL, program execution, and optional neural post-processing for refining results. It has been successfully applied across several areas of computer graphics. In 2D shape modeling, they are used in layout generation [53, 44], engineering sketch creation [15, 43, 48], and vector graphics synthesis by constructing programs that represent geometric shapes and their spatial relations [10, 16]. In 3D shape modeling, they facilitate inferring shape programs from existing 3D models [50, 32, 26, 27] or generating entirely new 3D shapes by training generative models on shape programs [56, 59, 60] or generate generate node graphs that define complex textures and materials [49, 20, 24] following the procedural modeling paradigm. Additionally, neuro-symbolic methods have been employed in human motion prediction [17, 37], reasoning [57, 33, 65] and generation [13, 40, 61], which leveraging visual-language foundation models to extract symbolic representations from visual data, facilitating complex activity reasoning by combining visual cues with symbolic logic [57].
Our approach aligns with the neurosymbolic paradigm by instruction-tuning LLMs to generate GarmentCode from various forms of design concepts. Similar to neurosymbolic models used in 2D/3D shape modeling and procedural texture generation, our method synthesizes structured, executable programs that define garment components, their relations, and parameters. This allows for precise, customizable generation of sewing patterns given multi-modal design inputs while using the symbolic power of GarmentCode to ensure geometric and structural accuracy.
@@ 3.2.2 Program Synthesis
As in Eq. 4, we use the finetuned DSL-GA $\Gamma_{ft}$ to encode the answers $\tau_{i}$ from MMUA and construct the condition input for $\Psi$ (we use an MLP to match the embedding dimension between $3,072$ in $\Gamma_{ft}$ and $128$ in $\Psi$). Notably, our token sequence length is fixed to the number of design parameters $|\mathcal{D}|=122$, regardless of the complexity of the pattern. This fixed-length representation is at least 10×10$\times$10 × compact than DressCode [21], whose sequence length is $1,500$ and scales with pattern complexity (see Sec. 4.1 for implementation details).
4 Experiments
4.1 Implementation Details
We use GPT-4V [6] for MMUA, and an instruction tuned version of Llama-3.2-3B for DSL-GA ($\Gamma$). The following sections contains the detailed explanation for the finetuned DSL-GA, and training details for the Projector $\Psi$.
4.1.1 Finetuning DSL-GA $\Gamma_{ft}$
To optimize the trade-off between computational cost and generation quality, we implemented the DSL generation agent (DSL-GA) using the Llama-3.2-3B-Instruct[51] model, fine-tuned over two epochs with LoRA (rank 16) and a learning rate of $5\times 10^{-4}$. All code generation experiments were conducted on a single NVIDIA GTX 4090. For multi-modal understanding tasks, GPT-4V was employed as the designated agent.
4.1.2 Training The Projector $\Psi$
The projector $\Psi$ is trained on the GarmentCodeData [31] dataset, which comprises approximately 115,000 garment samples draped on a standard A-pose body. We generate initial design descriptions for each sample using GPT-4V or rule-based inverse mapping from the ground truth design parameters for the sample, for example {minted}python if design.shirt.length.v ¿ 1.0: return ’shirt__length__long’ The token sequence length is fixed at $122$, which is equal to the number of design parameters in GarmentCode. The projection MLP and Transformer decoder are designed with feature dimensions of $128$. The MLP consists of $4$ intermediate layers, while the Transformer decoder includes $8$ layers. Training is conducted using the Adam optimizer with a learning rate of $5\times 10^{-4}$, a batch size of $16$, and completed on a single NVIDIA GTX 4090 within $10$ hours.
$L_{seq}=N_{p}\times(N_{e}\times L_{e}+\|R\|+\|T\|+N_{e}\times\|S\|)+2$
@@ 4.1.2 Training The Projector $\Psi$
Diversity # Panels $5.11_{\pm 2.76}$ $6.92_{\pm 9.63}$ $10.11_{\pm\textbf{19.53}}$ $\textbf{11.02}_{\pm 17.48}$
# Edges $5.48_{\pm 2.56}$ $6.84_{\pm 11.42}$ $5.79_{\pm 2.91}$ $6.24_{\pm 8.4}$
# Stitches $10.06_{\pm 10.51}$ $18.66_{\pm 74.58}$ $15.81_{\pm 34.97}$ $27.9_{\pm 96.61}$
Table 1: Quantitative comparisons between our method and SOTA sewing pattern generation methods in terms of generation quality and diversity. SSR is Simulation Success Rate, Agreement quantifies alignment with design prompts, and Aesthetic evaluates the visual appeal of the generated patterns. #Panels, #Stitches, and #Edges represent the average and variance (subscript) number of panels/stitches per pattern and edges per panel, respectively.
@@ 4.2 Quantitative Evaluation
Generation Quality is evaluated through three metrics: Simulation Success Rate (SSR), Agreement Score, and Aesthetic Score. The Simulation Success Rate (SSR) is calculated as the ratio of successfully simulated garments to the total number of generated sewing patterns, measuring the structural feasibility of the patterns. We prepared a dataset comprising 150 text prompts and 150 test images. For each sample, we generated sewing patterns using both our method and baseline methods, and simulated the patterns using GarmentCode’s simulation engine [30, 31] to compute the success rate. The Agreement and Aesthetic Scores were derived from a user study involving 30 professional pattern-makers. Each participant is asked to review 50 text and 50 image test samples. For each sample, we present the participants’ sewing patterns and simulated garments generated by our method and the baseline models, and ask them to rate each based on two criteria:
@@
For each criterion, participants could express a preference for either our method or the baseline or indicate that both methods were “comparable”. We then calculated the Agreement and Aesthetic Scores as the percentage of times each option was chosen over the total number of tested samples.
@@
Figure 7: Sewing Pattern Authoring with instructions. Starting from an original design, the system follows user instructions to adjust specific pattern elements. In Edited Design 1, the pants are modified to a skirt based on the command “CHANGE THE PANT TO SKIRT”. In Edited Design 2, the sleeves are shortened as requested. Finally, in Edited Design 3, the shirt is made sleeveless in response to the instructions. Note that, each modification accurately applies only to the specified parts, leaving the rest of the design unchanged.
@@ 6 Conclusion
Design2GarmentCode transforms multi-modal design concepts into precise sewing patterns using LMMs to synthesize parametric programs. It addresses challenges related to data requirements, computation, and the limited precision of neural network-based methods. The experimental results demonstrate the system’s ability to capture design details while maintaining structural integrity and geometric precision in generated patterns.
@@
Despite these advantages, Design2GarmentCode currently cannot substantially alter GarmentCode’s underlying structure and logistics, which impacts generation quality due to inherent limitations in GarmentCode’s design and modeling capabilities. For example, the range of upper garment patterns is limited, making it difficult to model personalized segmentations (Figure 10 (b)). Additionally, for designs like halter necks or strapless tops (Figure 10 (a)), GarmentCode cannot model fine straps, leading to potential simulation failures. These constraints restrict the system’s ability to accurately represent certain complex or customized garment designs.

### v3

@@ Abstract
Sewing patterns, the essential blueprints for fabric cutting and tailoring, act as a crucial bridge between design concepts and producible garments. However, existing uni-modal sewing pattern generation models struggle to effectively encode complex design concepts with a multi-modal nature and correlate them with vectorized sewing patterns that possess precise geometric structures and intricate sewing relations. In this work, we propose a novel sewing pattern generation approach Design2GarmentCode based on Large Multimodal Models (LMMs), to generate parametric pattern-making programs from multi-modal design concepts. LMM offers an intuitive interface for interpreting diverse design inputs, while pattern-making programs could serve as well-structured and semantically meaningful representations of sewing patterns, and act as a robust bridge connecting the cross-domain pattern-making knowledge embedded in LMMs with vectorized sewing patterns. Experimental results demonstrate that our method can flexibly handle various complex design expressions such as images, textual descriptions, designer sketches, or their combinations, and convert them into size-precise sewing patterns with correct stitches. Compared to previous methods, our approach significantly enhances training efficiency, generation quality, and authoring flexibility. Project page: https://style3d.github.io/design2garmentcode.
@@ 2.2 LLMs for Program Synthesis
Recent advancements in program synthesis and code generation using large language models (LLMs) have laid essential groundwork for systems like Design2GarmentCode, which generate structured garment code from multi-modal design inputs. Earlier researches like Codex [11] and AlphaCode [36] demonstrated the effectiveness of LLMs in generating complex, task-specific code with high syntax accuracy, showcasing potential in scenarios requiring precise parametric coding. These models [43, 59] highlight how LLMs, when sufficiently trained, can transform natural language inputs into executable code, a capability directly relevant to generating garment codes that follow complex pattern-making syntax.  [55, 19, 9, 20, 63] explore multi-modal models that integrate visual aids, such as flowcharts and UML diagrams, into LLM training to enhance models’ comprehension of complex structures and flow. These models particularly emphasize the need for semantic understanding and adaptability, which are critical in working with domain-specific languages (DSLs) like GarmentCode.
@@ 2.3 Neurosymbolic Models
Procedural/symbolic models and learned/neural models have complementary strengths and weaknesses. Neurosymbolic models [47] tend to combine the strengths of both paradigms and propose to generate visual data using symbolic programs augmented with AI/ML techniques. The neurosymbolic pipeline typically includes task specification, program synthesis using a DSL, program execution, and optional neural post-processing for refining results. It has been successfully applied across several areas of computer graphics. In 2D shape modeling, they are used in layout generation [54, 45], engineering sketch creation [16, 44, 49], and vector graphics synthesis by constructing programs that represent geometric shapes and their spatial relations [10, 17]. In 3D shape modeling, they facilitate inferring shape programs from existing 3D models [51, 33, 27, 28] or generating entirely new 3D shapes by training generative models on shape programs [57, 60, 61] or generate generate node graphs that define complex textures and materials [50, 21, 25] following the procedural modeling paradigm. Additionally, neuro-symbolic methods have been employed in human motion prediction [18, 38], reasoning [58, 34, 66] and generation [13, 41, 62], which leveraging visual-language foundation models to extract symbolic representations from visual data, facilitating complex activity reasoning by combining visual cues with symbolic logic [58]. Our method leverages a neurosymbolic approach by instruction-tuning LLMs to generate GarmentCode design configurations and component programs from diverse design concepts, ensuring geometric and structural accuracy.
@@ 3.2.2 Program Synthesis
As in Eq. 4, we use the finetuned DSL-GA $\Gamma_{ft}$ to encode the answers $\tau_{i}$ from MMUA and construct the condition input for $\Psi$ (we use an MLP to match the embedding dimension between $3,072$ in $\Gamma_{ft}$ and $128$ in $\Psi$). Notably, our token sequence length is fixed to the number of design parameters $|\mathcal{D}|=122$, regardless of the complexity of the pattern. This fixed-length representation is at least 10×10$\times$10 × compact than DressCode [22], whose sequence length is $1,500$ and scales with pattern complexity (see Sec. 8 for implementation details).
@@
CLIPScore 0.2386 0.2438 / /
F-Score 0.616 / 0.708 0.829
CD 15.77 / 9.7 2.091
Diversity # Panels $5.11_{\pm 1.66}$ $6.92_{\pm 3.10}$ $10.11_{\pm\textbf{4.42}}$ $\textbf{11.02}_{\pm 4.18}$
# Edges $5.48_{\pm 1.60}$ $6.84_{\pm 3.38}$ $5.79_{\pm 1.71}$ $6.24_{\pm 2.90}$
# Stitches $10.06_{\pm 3.24}$ $18.66_{\pm 8.64}$ $15.81_{\pm 5.91}$ $27.9_{\pm 9.83}$
Table 1: Quantitative comparison of our method against state-of-the-art (SOTA) sewing pattern generation techniques in terms of quality and diversity. SSR (Simulation Success Rate) indicates the feasibility of simulated garment assembly, while Agreement measures alignment with design prompts, and Aesthetic evaluates the visual preference of the generated patterns. CLIPScore assesses text-image consistency, whereas Chamfer Distance (CD) and F-Score quantify geometric accuracy. #Panels, #Stitches, and #Edges denote the mean and standard deviation (subscript) of the number of panels, stitches per pattern, and edges per panel.
4 Experiments
@@ 4.1 Quantitative Evaluation
Generation Quality is evaluated through three metrics: Simulation Success Rate (SSR), Agreement Score, and Aesthetic Score. The Simulation Success Rate (SSR) is calculated as the ratio of successfully simulated garments to the total number of generated sewing patterns, measuring the structural feasibility of the patterns. We prepared a dataset comprising 150 text prompts and 150 test images, covering a wide variety including tops (78), pants (76), skirts (38), dresses (80), and suits (28). For each sample, we generated sewing patterns using both our method and baseline methods, and simulated the patterns using GarmentCode’s simulation engine [31, 32] to compute the success rate. The Agreement and Aesthetic Scores were derived from a user study involving 30 professional pattern-makers. Each participant is asked to review 50 text and 50 image samples generated by our and the baseline models, and assess their preference based on sewing patterns and simulated garments according to:
@@
For each criterion, participants could express a preference for either our method or the baseline or indicate that both methods were “comparable”. We calculate the Agreement and Aesthetic Scores as the percentage of times each option was chosen over the total number of tested samples.
@@
Figure 7: Sewing Pattern Authoring with instructions. Starting from an original design, the system follows user instructions to adjust specific pattern elements. In Edited Design 1, the pants are modified to a skirt based on the command “CHANGE THE PANT TO SKIRT". In Edited Design 2, the sleeves are shortened as requested. Finally, in Edited Design 3, the shirt is made sleeveless in response to the instructions. Note that, each modification accurately applies only to the specified parts, leaving the rest of the design unchanged.
@@ 6 Conclusion
Design2GarmentCode transforms multi-modal design concepts into precise sewing patterns using LMMs to synthesize parametric programs. It addresses challenges related to data requirements, computation, and the limited precision of neural network-based methods. The experimental results demonstrate the system’s ability to capture design details while maintaining structural integrity and geometric precision in generated patterns. Despite these advantages, Design2GarmentCode currently cannot substantially alter GarmentCode’s underlying structure and logistics, which impacts generation quality due to inherent limitations in GarmentCode’s design and modeling capabilities (Supp. 10).
Acknowledgments
This work was supported by Key R&D Program of Hangzhou (No.2024SDZ1A20). We sincerely thank Fuqi Wang for his valuable support in validation data collection, material modeling and animation.
Supplementary Material
7 Validations in Design2GarmentCode
7.1 Rule-based Validation
Rule-based validation primarily addresses issues of completeness and hallucination during the MMUA’s generation process. With prompts generated by the DSL-GA containing over 100 questions, the MMUA often struggles to provide comprehensive answers in a single attempt. Additionally, due to the inherent hallucination tendencies of LLMs, some responses may fall outside the reasonable parameter range defined by GarmentCode. To mitigate these issues, we compare the MMUA’s responses against a predefined complete question space to verify whether all questions have been adequately addressed before program synthesis. Each response is further validated to ensure it falls within GarmentCode’s permissible parameter space. Questions with either missing or invalid answers are sent back to the MMUA for re-evaluation, with a maximum of two validation loops to refine the outputs.
7.2 MMUA Design Comparison
During design comparison we ask the MMUA to compare the output design image versus the design input and propose modification suggestions to DSL-GA to edit the generated pattern. Design comparison is especially useful for image-guided generation, where the output design image is rendered from the draped garment mesh under a similar viewpoint to the input image, we use TokenHMR [15] to estimate the camera pose and rough human pose from the input design image. The prompt used for design comparison is given in Figure 10.
Figure 10: Prompt for MMUA during design comparison.
Figure 11: LMM-based interface for Design2GarmentCode, built upon the original GarmentCode GUI. The chat interface (left) allows users to provide natural language design descriptions or upload reference images or sketches, facilitating multi-modal design parsing into executable GarmentCode programs. We use the original GarmentCode execution engine to turn the generated program into 3D garments.
Figure 12: User study interface for evaluating sewing pattern generation quality. For each test input (Original image on the left for image-based evaluation or Text Description on the right for text-based evaluation), participants are presented with the simulation results of sewing patterns generated by Design2GarmentCode and a baseline method. Users are asked to evaluate the results based on two criteria: agreement with the input description and overall sewing pattern quality. If unsure, participants can select the "Unsure" option.
8 Implementation Details
We use GPT-4V [6] for MMUA, and an instruction tuned version of Llama-3.2-3B for DSL-GA ($\Gamma$). The following sections contains the detailed explanation for the finetuned DSL-GA (Supp. 8.1), and training details for the Projector $\Psi$ (Supp. 8.2).
8.1 Finetuning DSL-GA $\Gamma_{ft}$
To optimize the trade-off between computational cost and generation quality, we use Llama-3.2-3B-Instruct[52] as the base model for DSL-GA, fine-tuned over two epochs with LoRA (rank 16) and a learning rate of $5\times 10^{-4}$ on a dataset with 583 hierarchically defined NL-DSL pairs from GarmentCode’s public code repository. All code generation experiments were conducted on a single NVIDIA GTX 4090. For multi-modal understanding tasks, GPT-4V was employed as the designated agent.
8.2 Training The Projector $\Psi$
The projector $\Psi$ is trained on the GarmentCodeData [32] dataset, which comprises approximately 115,000 garment samples draped on a standard A-pose body. We generate initial design descriptions for each sample using GPT-4V or rule-based inverse mapping from the ground truth design parameters for the sample, for example {minted}python if design.shirt.length.v > 1.0: return ’shirt__length__long’ The token sequence length is fixed at $122$, which is equal to the number of design parameters in GarmentCode. The projection MLP and Transformer decoder are designed with feature dimensions of $128$. The MLP consists of $4$ intermediate layers, while the Transformer decoder includes $8$ layers. Training is conducted using the Adam optimizer with a learning rate of $5\times 10^{-4}$, a batch size of $16$, and completed on a single NVIDIA GTX 4090 within $10$ hours.
$L_{seq}=N_{p}\times(N_{e}\times L_{e}+\|R\|+\|T\|+N_{e}\times\|S\|)+2$
8.3 Inference Interface
For more convenient inference, we build an intelligent chat-based interface integrated into the original GarmentCode [31] GUI (Figure 11). The chat interface (left panel) enables users to provide natural language descriptions, upload reference images, or supply design sketches, facilitating multi-modal design parsing into GarmentCode-compliant programs which are then passed to the GarmentCode execution engine. The engine generates sewing patterns and 3D garment simulations (right panel). This interactive interface provides an intuitive environment for creating, editing, and refining sewing patterns, significantly improving accessibility for users without extensive expertise in parametric pattern-making. We provide a recording to demostrate the inference process in demo.mp4.
Figure 13: Example answers from Llama 3.2 3B when prompted with “How to draft a basic upper body bodice?”.
@@ 8.3 Inference Interface
8.4 User Study Interface
To evaluate the quality of sewing pattern generation, we design a user study interface tailored for comparison (Figure 12). For each test input—either an original image (for image-based evaluation) or a text description (for text-based evaluation)—the interface presents participants with simulated garment results generated by Design2GarmentCode and a baseline method. Participants assess the results based on two criteria: Agreement, which measures how well the generated patterns align with the input description, and Aesthetic, which evaluates the structural integrity and aesthetic appeal quality of the patterns. An "Unsure" option is available for cases where a clear preference cannot be determined, ensuring unbiased and flexible feedback.
9 LMM Prompting Details
9.1 Pattern Drafting Test
As outlined in Sec. 3.2.1, a key prerequisite for Design2GarmentCode is the presence of embedded pattern-drafting knowledge in pre-trained large models. To assess this capability, we prompted models like O1-preview and LLama 3.2 3B Instruct with the question, "How to draft a basic upper body bodice?". These models produced step-by-step drafting instructions in natural language, including commands such as: "STEP 1: Take Your Measurements," and "STEP 2: Draw the Center Front Line, Draw the Shoulder Line, Draw the Armhole, Draw the Side Seam (Measure the distance from the underbust measurement and divide it by 4. Mark this distance from the armhole point down to the waist. Draw a vertical line to represent the side seam)." Figure 13 showcases sample outputs from Llama-3.2-3B Instruct which we used as baseline for DSL-GA.
9.2 Prompting for MMUA
Based on different input design modalities and tasks, we assigned five specific tasks to the MMUA.
Task 1: Identify the image, extract answers for each prompt question based on the image, and combine them to form the recognized clothing information. This task establishes the relationships between parameters and the questions corresponding to each parameter. It serves as the foundation for all subsequent tasks.
Task 2: Generate clothing information based on text. Building on Task 1, this task generates clothing prototype information according to user preferences.
Task 3: Retrieve existing clothing information and modify the clothing design according to the user’s ideas.
Task 4: Input stress test images along with the current clothing information from the text space. MMUA interprets the colors in the image as stress levels—red, yellow, or similar colors indicate areas that are too tight. MMUA dynamically adjusts the clothing information to reduce stress.
Task 5: Compare previously generated clothing simulation images, their corresponding clothing information, and the original input image. Identify differences between the simulation and the original image, and dynamically adjust the clothing information to make the final simulation image more closely resemble the original.
Figure 15: Example of original design configurations with numerical values and modified design configurations with only selective parameters.Example of original design configurations with numerical values and modified design configurations with only selective parameters.
As discussed in Sec. 3.2.2, due to the probabilistic nature of LMMs, the MMUA struggles to accurately estimate numerical values in the design configuration. Therefore, we limit the MMUA’s task to answering multiple-choice questions, with responses formatted as a selective parameter list. Fig. 15 illustrates example parameters before (design_cfg_num) and after (design_cfg_slc) modification. The complete prompt will be publicly available with Design2GarmentCode code base.
10 Limitations
A limitation of Design2GarmentCode is its current inability to substantially modify GarmentCode’s underlying structure and logic, which impacts the generation quality due to inherent constraints in GarmentCode’s design and modeling capabilities. For example, the range of upper garment patterns is limited, making it difficult to model personalized segmentations (Figure 14 (b)). Additionally, for designs like halter necks or strapless tops (Figure 14 (a)), GarmentCode cannot model fine straps, leading to potential simulation failures. These constraints restrict the system’s ability to accurately represent certain complex or customized garment designs.
11 Additional Results
In the following, we present additional experimental results in text-, image-, and sketch-guided pattern generation and highlight the versatility and effectiveness of our approach across various modalities.
Figure 16: Additional Text-guided generation results. From left to right: input text; output from MMUA; generated sewing pattern; simulation results.
Figure 17: Additional Image-guided generation results. From left to right: input image; output from MMUA; generated sewing pattern; simulation results.
Figure 18: Additional Sketch-guided generation results. From left to right: input sketch; output from MMUA; generated sewing pattern; simulation results.
Figure 19: Additional sewing pattern editing results.Starting from the original sewing pattern on the far left, the system applies user instructions to edit the pattern. The left side of each arrow represents the original pattern, while the right side displays the edited result.